Recent Log Entries
4th Pass Separation Check
Crew Chief Shift Log, 2023-02-14 SWING
Log Number: 4131321Logbook(s): Tags: Body:
CEBAF SHIFT LOG
Shift Summary
Start of shift:
GP-PL-GP
Crew Chief Timesheet Summary: 14 Feb 2023 SWING
Log Number: 4131317Logbook(s): Tags: Body:
Accelerator
PHYSICS
STUDIES
RESTORE
ACC
DOWN
OFF (SAD)
Planned
8
0
0
dogCalc V0-12 download
Downtime Incident CLOSED: 1L12 Dropped Out of HV
Phaser Results
The OPS DRVHs for the following cavities have been updated: 1L11-2
Log Number: 4131304Logbook(s): Tags: Body:
The OPS DRVHs for the following cavities have been updated:
1L11-2
Username: anthony
Date/Time: 02-14-23_22:41:46
RF Status: Good
Tuner Status: Good
Comment: DRVH=7.400 Cavity SOSing above this level.
Previous OPS DRVH: 8.2
Special Instructions:
HYST AREA 6-2-4 Cycles Magnets for LEM
LEM North Linac
FudgeIt
FSD Masking Tool: Successful System Masking for: BeamLossMonitor
Log Number: 4131300Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: anthony
Areas: All
BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T0
User: anthony
Areas: All
BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T0
HYST AREA 6-2-4 Cycles Magnets for LEM
HYST AREA 6-2-4 Cycles Magnets for LEM
FSD Masking Tool: Successful Destination masking to: HallA,HallB,HallC,HallD
LEM South Linac
LEM North Linac
1L12 HV and cavity 4 No KCCU
Log Number: 4131287Logbook(s): Entry Makers: kernsBody:
1L12 not going to HV upon beam permit. Found CPS fan off and 3 ac fuses keeping it out of HV enable. Closed all back up and in HV and half an hour later cav. 4 kly. resistor blew and no KCCU. Replaced resistor and restored zone to RF. Resistor and RF restore by RF guru Larry the RF Man and thanks to Ben for getting substance (pizza) to give us the needed energy to get another shift in and beam going. L. Farrish and C. Kerns.